OpenStudy (anonymous):

A wall clock is 12 inches in diameter. The clock has twelve equally spaced numbers. What is the distance around the edge of the clock from 3 to 6?

OpenStudy (rulnick):

pi d / 4 = pi (12 in) / 4 = 3pi inches

OpenStudy (rulnick):

... so about 9.425 inches.